OpenStudy (anonymous):

find the intergral of 1+cosAcosx/cosA+cosx

OpenStudy (anonymous):

is it\[\int \frac{1+\cos A \cos x}{\cos A +\cos x} \text{d}x\]? and A is a constant?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i forgot to mention that

OpenStudy (anonymous):

maybe weierstrass substitution\[t=\tan \frac{x}{2}\]

OpenStudy (anonymous):

what is weierstrass substitution?!?!?!?!

OpenStudy (anonymous):

\[xcos(a)+\sin(a)\log(\cos(a \div 2 - x \div 2)) - \sin(a)\log(\cos(a \div 2 - x \div 2))\]

OpenStudy (anonymous):

"weierstrass

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[a=\cos A\]\[ \frac{1+a \cos x}{a +\cos x}=\frac{1+a\cos x+a^2-a^2}{a +\cos x}=a+\frac{1-a^2}{a +\cos x} \]so\[I=\int (a+\frac{1-a^2}{a +\cos x})\text{d}x=ax+(1-a^2) \int \frac{1}{a+\cos x }\text{d}x\]

OpenStudy (anonymous):

best method for later integral is weierstrass substitution

OpenStudy (anonymous):

how do u know when to put a^2(+ &-)?

OpenStudy (anonymous):

i just wanted to simplify it a bit so i looked at denum and i thought an a^2 inside of a cos x in num will make my work easier

OpenStudy (anonymous):

your damn good at inte......

OpenStudy (anonymous):

:)